OpenStudy (anonymous):

Rectangle A has an area of 4-x^2. Rectangle B has an area of x^2+2x-8. In simplest form, what is the ratio of the area of Rectangle A to the area of Rectangle B? Show your work.

OpenStudy (studyhard):

I could only get this far, hope it helps! 4-x^2=a^2-b^2=(a+b)(a-b)=(2-x)(2+x). For x^2+2x-8, you must cross multiply to get (x+4)(x-2). Sorry I couldn't get any further!

OpenStudy (studyhard):